Joan saved $21,000 in two banks. One bank gives 4 3/5% annual interest, and the other gives 5 1/2% annual interest. How much did she save in each bank if she received a combined annual interest rate of 5% on her savings?
where to even begin?

Answer by stanbon(75887)

Joan saved $21,000 in two banks. One bank gives 4 3/5% annual interest, and the other gives 5 1/2% annual interest. How much did she save in each bank if she received a combined annual interest rate of 5% on her savings?
where to even begin?
----
Using one variable:
Interest + Interest = Interest
----------
0.046x + 0.055(21000-x) = 0.05*21000
----
Multiply thru by 1000 to get:
46x + 55*21000 - 55x = 50*21000
-9x = -5*21000
----
x = (5/9)21000
x = $11,666.67 (amt. invested at 4 3/5%)
---
21000-x = $9333.33 (amt. invested at 5 1/2%
============================================
Cheers,
Stan H.
===============